# Artificial Neural Networks for Magnetoencephalography: A review of an emerging field

### v1

@@ 3.2 Classification studies
As explained in section 2, we split the studies in this category into three MEG-based classification problems: Decoding problems, Clinical studies, and Brain-Computer Interfaces (BCI).
Decoding studies: Generally speaking, decoding studies aim better to understand the neural correlates of behavior and cognitive processes. This is typically accomplished by using ML or statistical inference. This approach determines the neural responses that underlie specific sensory, motor, or higher-order cognitive functions by revealing the features that exhibit the strongest separation between classes. In this category, 14 studies across four groups of researchers use ANNs for speech decoding or speech-related BCI ([29, 46, 51, 50, 23, 21, 48, 49, 22, 24, 18, 58, 38]). Three studies specifically focus on using ANNs for auditory stimuli decoding ([36, 20, 34]). Additionally, 10 studies use them for visual stimuli decoding ([25, 26, 27, 28, 31, 33, 34, 42, 44, 40]). The remaining eight decoding studies use ANNs for affect states decoding ([30]), age decoding ([43]), hand gestures ([37, 35]), rhythm decoding ([19]), text decoding ([41, 45]), or decoding methods ([32]).
Clinical studies: The general goal of studies classified as Clinical is to predict or detect diseases or anomalies in patients. Thirteen studies use ANNs for clinical purposes or computer-aided diagnosis (CAD). Two focus on epilepsy-related signal properties detection ([71, 66]). Two studies focus on Alzheimer’s disease early detection [62] or evolution [68]. Three focus on neurological disease diagnostic ([60, 70, 69]), and the remaining studies focus on schizophrenia detection [67], autism detection in children population [72], PTSD severity evaluation [65], migraine diagnostic [61], Mild traumatic brain injury detection [63] or depression and bipolar disorder detection [64]. Although not classified as clinical studies, among the 15 studies sub-classified as ’Event Detection’, 13 are focused on detecting and/or locating epileptic spikes ([78, 77, 84, 85, 82, 79, 80, 83, 74, 76, 73, 75, 86]). Therefore a combined total of 14 studies out of the 69 classification studies are epilepsy-related.
@@
BCI studies: The third subcategory regroups the BCI studies, which generally assess the feasibility of building brain-controlled devices that rely on predicting the subject’s intentions from their neural data. BCI studies cover a wide range of approaches. Still, most consist either of offline assessments of previously recorded data or online examination of intention decoding in a closed-loop setting. BCI is one of the neuroscience domains in which ML made its earliest incursions (e.g. decoding arm movement direction and kinematics from neuronal activity in non-human primates). Of the remaining six studies subcategorized as BCI, five use ANNs for motor tasks ([52, 53, 56, 55, 54]), and finally, in [57], the authors investigate how ANNs can improve the state-of-the-art of BCI.
Standard ML algorithms (e.g. support-vector machine, linear discriminant analysis, or random forest) are increasingly used in MEG-based decoding studies. However, according to our survey, the first ANN-based MEG classification studies were published in a journal only in 2016 [30]. To date, the MEG studies that have applied ANNs for classification purposes are numerous and very different from one another. In particular, due to differences in the goal and rationale, the 69 classification studies vary substantially in terms of number of subjects (ranging from 2 [29] to 646 [43]). However, most studies, 21 of the 56 that mentioned it, use more than 1000 samples in their dataset, as seen in Fig. 1. Thirteen studies in this category did not mention the number of samples. Seventeen papers in this category have their codes publicly available or available upon request. Ten of these studies were published in 2023 or later. We will now provide an overview of the datasets used, followed by an overview of the network architectures implemented by the various papers.
3.2.1 The data
The studies in this group used MEG data sampled at frequencies between 50Hz and 2400Hz. Still, nine studies omitted details about the sampling frequency, in most cases because the authors extracted features directly from the raw signals, making the original sampling rates less pertinent([61, 62, 72, 71, 30, 114, 115, 53, 19]).
Typically, across all studies in this category, the MEG data were preprocessed before being fed as input to a neural network. This preprocessing phase generally includes down-sampling, band-pass filtering, and either de-noising and/or artifact removal. Of the 69 studies in this category, 12 studies (Decoding: [21, 22, 23, 29] BCI: [46, 47, 48, 49, 51, 53], Clinical: [64], Event Detection: [85]) use Discrete Wavelet Transform to increase Signal-to-Noise Ratio (SNR) in the data as a preprocessing step. Notably, these 12 studies are spread across only four distinct research groups representing a small portion of the compendium. Five studies employ alternative de-noising techniques ([28, 42, 69, 70, 77]). Additionally, 12 studies use ICA to remove heart and/or eye movement artefacts ([34, 37, 44, 57, 72, 67, 78, 81, 69, 70, 35, 40]). Three studies use Fieldtrip artefact removal algorithm ([68, 61, 62]) or EEGlab eye artefact removal ([18]). The remaining 38 studies do not explicitly mention de-noising or artifact removal techniques in their preprocessing pipeline.
@@ 3.2.1 The data
Among all classification papers, 27 studies make use of hand-crafted features. While most of these studies train the network uniquely with hand-crafted features ([53, 61, 46, 47, 66, 51, 62, 63, 68, 75, 73, 73, 71, 23, 30, 22, 24, 29, 33, 64, 70, 59, 85]), a few use a combination of hand-crafted features and raw MEG data ([65, 20, 60, 69]. Among the reviewed studies, spectral power and functional connectivity matrices were the two most frequently used types of hand-crafted features.
Data augmentation is often necessary when using ANNs, especially considering that small datasets can hinder generalization performances and that acquiring MEG data is expensive. Among the surveyed classification studies, 12 make use of augmentation techniques ([76, 46, 23, 18, 64, 39, 42, 40, 70, 38, 45, 86]). These studies use some form of linear shifting and/or sensor shuffling or mix-up regularization ([136]). An example of linear shifting consists of selecting a different cropping of a trial or segment around the cue in the temporal dimension. The shifted samples end up having the same length but different onsets.
@@
In the studies that used the original MEG data as input rather than hand-crafted features, the reported length of the segments used as samples for training and testing the ANNs varied between 140 and 2,400. Unfortunately, a few papers do not mention the length of the segments used as input for their neural network ([26, 31, 19, 69]).
3.2.2 The networks
The most commonly used ANNs for classification-type studies are Convolutional Neural Networks (CNNs). The second most used architecture was Multi-Layer Perceptrons (MLP), followed by Recurrent Neural Network (RNN) architectures. For the CNNs, the depth of the networks varied between one and 144 hidden layers. The papers that used MLPs essentially limit themselves to six hidden layers.
The CNNs were designed to mimic the human visual cortex. It comprises successive convolution layers that can learn abstract representations of the data. The convolution layers are generally followed by two to three Fully Connected (FC) layers that will perform the final classification. Long Short-Term Memory Networks (LSTMs) are an advanced form of Recurrent Neural Networks (RNNs), addressing some issues with long-term dependencies in the original RNN design. RNNs can learn sequential/temporal relationships in data and predict the next element of a sequence. Autoencoders are MLP architectures where the first half of the networks consists of Fully Connected (FC) layers of decreasing sizes, followed by FC layers of increasing sizes up to the size of the input. Such architecture aims to learn an ideally lossless compressed representation of the data. Sparse Autoencoders (SAEs) are a special case of AEs where a sparsity constraint is added to the hidden units: most units composing the hidden FC layers will randomly be set to zero for any given input. More info on these networks can be found in [1].
Like any other parametric algorithm, ANNs have parameters that significantly influence their ability to learn and acquire a good data representation. These parameters define the architecture’s depth and shape, and the optimal choices vary depending on the application. This section provides an overview of the predominant architectures and parameter configurations employed in the studies categorized under ’classification’.
@@ 3.2.2 The networks
Most of the CNN architectures used in the reported MEG classification papers comprise the input layer followed by multiple blocks convolution layers with maxpool and/or Rectified Linear Unit (ReLU) layers. At the end, one or two fully connected layers are followed by an activation function. For the convolution layers, the critical parameters are filter size, stride, padding, and number of channels. For fully connected layers, the crucial parameter is the number of units, also called the width of the layer.
@@
Beyond the architecture used and data (pre-)processing, details about training parameters, loss function, optimizer, and regularization techniques are essential to replicate the results of such studies. While training parameters may not have as significant an impact as architecture parameters, they remain essential and can heavily impact training time and performance. Out of the 69 classification studies, 17 do not mention training parameters. The remaining studies generally mention learning rate, batch size, loss function, optimization method, and regularization techniques. Batch sizes, regularization techniques (dropout, L1 and L2 norm weight usage), number of epochs, early stop criteria, learning rates, and momentum parameters are crucial to watch out for. However, providing a detailed description for each parameter across all surveyed articles would not be very useful, as each study has its own setting, and no general rule can be extracted. However, further details on these parameters can be found in the complete table (https://tinyurl.com/ub3s5mr).
@@
Among the studies that mentioned the optimization algorithm used for training their network, 29 studies used Adam [137]) or some variation thereof, while four used Stochastic Gradient Descent (SGD). Adam, short for Adaptive Moment Estimation, uses both the first-order and second-order moments of the gradient to adapt the learning rate, while SGD only uses the first-order moment. Adam typically converges faster than SGD.
@@
Validation techniques are essential to training any model in a classification setting, enabling the proper evaluation of the model’s performance. In a setting where generalizing to new subjects is essential, such as BCI or CAD, it is capital to exclude multiple subjects’ data from the training process. This ensures the test of the trained algorithm’s generalization capabilities on entirely new and unseen data. It is also essential to evaluate and fine-tune the algorithm’s performances on training and validation datasets that do not share information with each other or the test set. Among the 69 studies in this category, 26 used the K-fold validation technique (K varying from two to ten), 28 used a simple train/valid/test set split, 10 used Leave One Out or Leave P out at the subject level. Five did not mention any validation technique in their analysis pipeline. The train/valid/test random splitting method is generally preferred in deep learning settings where the amount of data and the network size make training times very long, sometimes spanning multiple days. A large amount of data will dilute the split’s impact on the algorithm’s final performance. Note that when data is scarce, it is better to use bootstrapping or cross-validation to validate the model’s performance before evaluation on a test set, as it will lead to a less noisy estimate of performances ([138]). The Leave One Subject Out (LOSO) cross-validation strategy is preferred when subject generalization is essential, and the amount of available data is small. Note that when using K-fold or random split, it is crucial to think about whether it is important or not that the algorithm has seen parts of one subject’s data in the training set before validating or testing. Otherwise, the algorithm might end up identifying the subjects instead of solving the task it was designed for.
@@
ANNs are powerful and can tackle a large set of problems; however, they present challenges in terms of interpretation, implementation, and tuning. In some instances, it is unnecessary to use ANNs for problems that simpler algorithms can solve. This is why it is essential when using ANNs to check whether we are using the right tool for the right problem by computing a baseline of performances that can be reached with more classical machine-learning approaches. Unlike ANNs, proper feature selection is essential for classical algorithms, particularly when handling high-dimensional like MEG data. Among the papers in the Classification category, 26 out of the 69 studies did not compare their ANN’s performance to simpler algorithms’ performances.
@@
A critical part of decoding studies is understanding what aspect of the data allowed decoding. It is essential to find where and how information is encoded in the data to understand more about the brain. When using ML algorithms, especially ANNs, it can be hard to make sense of what is happening in the latent space. This is why using visualization tools will help with the interpretability of the network. Of the 69 papers in the classification category, only 17 include some interpretation or visualization techniques applied to the network. Only one of these study is a BCI sub-categorized study ([57]), two are event detection studies ([78, 84]), four are clinical sub-categorized studies ([68, 61, 64, 72]), and the remaining 10 are decoding studies ([26, 27, 34, 42, 32, 18, 44, 40, 41, 35]).

### v3

@@ 3.2 Classification studies
3.2.1 Study aims and subcategories
As explained in section 2, studies in the ’Classification’ category were divided into four main subcategories based on their primary goal: BCI, decoding problems, clinical applications, and event detection.
@@ 3.2.1 Study aims and subcategories
BCI studies: BCI studies assess the feasibility of building brain-controlled devices that rely on predicting the subject’s intentions from their neural data. BCI studies cover a wide range of approaches. Still, most consist either of offline assessments of previously recorded data or online examination of intention decoding in a closed-loop setting. BCI is one of the neuroscience domains in which ML made its earliest incursions (e.g. decoding arm movement direction and kinematics from neuronal activity in non-human primates). Of the 13 studies subcategorized as BCI, five use ANNs for motor tasks ([94, 93, 99, 102, 103]), seven focus on speech-related BCI ([91, 92, 96, 97, 98, 100, 101]), and finally, in [95], the authors investigate how ANNs can improve the state-of-the-art of BCI.
Decoding studies: Generally speaking, decoding studies aim to better understand the neural correlates of behavior and cognitive processes. This is typically accomplished by using ML or statistical inference. This approach determines the neural responses that underlie specific sensory, motor, or higher-order cognitive functions by revealing the features that exhibit the strongest separation between classes. In this category, eight studies use ANNs for speech decoding ([50, 53, 57, 61, 60, 56, 101, 72]). Three studies specifically focus on using ANNs for auditory stimuli decoding ([65, 63, 54]). Additionally, 10 studies use them for visual stimuli decoding ([52, 51, 55, 62, 58, 70, 54, 69, 67, 68]). The remaining decoding studies use ANNs for affect states decoding ([49]), age decoding ([66]), hand gestures ([71, 74]), rhythm decoding ([64]) or text decoding ([73, 75, 77]).
Clinical studies: we describe clinical studies, which aim to predict or detect diseases or anomalies in patients. Thirteen studies use ANNs for clinical purposes or CAD. Two focus on epilepsy-related signal properties detection ([87, 80]). Two studies focus on Alzheimer’s disease early detection ([83]) or evolution ([82]). Three focus on neurological disease diagnostic ([79, 90, 89]), and the remaining studies focus on schizophrenia detection ([84]), autism detection in children population ([88]), PTSD severity evaluation ([81]), migraine diagnostic ([78]), mild traumatic brain injury detection ([85]) or depression and bipolar disorder detection ([86]).
Event detection studies: Lastly, this subcategory encompasses research using ANNs to identify specific, often transient, events within continuous or segmented MEG recordings. A dominant application within the reviewed literature involves the automatic detection and sometimes localization of pathological neural events, particularly epileptic spikes or high-frequency oscillations, which are crucial biomarkers for epilepsy diagnosis and treatment planning ([105, 165, 114, 116, 104, 111, 106, 113, 110, 108, 112, 109, 117]). Other studies target different types of events, such as specific visual targets within rapid presentations ([107]). Collectively, event detection studies contributed significantly to the ’Classification’ category, with epilepsy-related applications being particularly numerous (15 studies across Clinical and Event Detection subcategories combined).
3.2.2 The pipeline
Across all classification studies, the typical pipeline involves using MEG recordings as input to ANNs with the goal of predicting class labels associated with sensory, motor, cognitive, or clinical states. While the precise implementation varies, most studies follow a common structure that includes preprocessing the MEG signals, optionally extracting features, and feeding the resulting data to a classifier—most often a CNN. The input of the model may consist of hand-crafted features, raw MEG signals, or a combination of both. These inputs are typically segmented into trials or epochs of fixed duration, which are then used to train and evaluate the model. The network outputs a predicted class label or probability distribution over target categories.
Despite this shared pipeline, classification goals, model inputs, and preprocessing strategies differ substantially depending on the application, and no standard protocol has emerged across the 70 studies. Figure 4 (left panel) illustrates a schematic representation of this typical Classification pipeline, alongside comparable diagrams for the Modeling and Other application categories.
Figure 4: Representative workflows for applying Artificial Neural Networks (ANNs) to MEG data across the main application categories identified in this review. (Left) ’Classification’ pipeline: Typically involves preprocessing MEG data, feeding it to an ANN classifier, and evaluating prediction accuracy. (Center) ’Modeling’ pipeline: Often compares representations (activations) extracted from ANNs processing stimuli with corresponding preprocessed MEG data, using similarity analysis (e.g., RSA, neural predictivity) and evaluating against baselines. (Right) ’Other’ pipeline: Encompasses methodological applications like preprocessing or source localization using ANNs, often evaluated by comparison with classical approaches or metrics like AUC or spatial error.
3.2.3 The data and preprocessing
Participant numbers and dataset sizes varied considerably across the 70 reviewed classification studies. Due to differences in the goal and rationale, the number of subjects varied substantially (ranging from 2 [50] to 646 [66]). Sample sizes (trials or epochs) also varied widely; while a majority of studies reporting this information (21 out of 56) used more than 1000 samples, thirteen studies did not specify this number (see figure 2 for the distribution).
@@ 3.2.3 The data and preprocessing
Among all classification papers, 27 studies make use of hand-crafted features. While most of these studies train the network exclusively with hand-crafted features ([93, 78, 91, 97, 80, 92, 83, 85, 82, 109, 112, 112, 87, 53, 49, 61, 60, 50, 70, 86, 90, 115, 116]), a few use a combination of hand-crafted features and raw MEG data ([81, 63, 79, 89]). Among the reviewed studies, spectral power and functional connectivity matrices were the two most frequently used types of hand-crafted features.
@@
In the studies that used the original MEG data as input rather than hand-crafted features, the reported length of the segments used as samples for training and testing the ANNs varied between 125 and 2,400. This corresponds to trial durations ranging from approximately 41.67 ms to 10,000 ms, with a median of 1,063 ms across the studies that reported this information. Unfortunately, a few papers do not mention the length of the segments used as input for their neural network ([51, 58, 64, 89]).
Data augmentation techniques are generally recommended in deep learning to improve model generalization and mitigate overfitting ([1]), but was however employed relatively infrequently in the surveyed classification studies. Specifically, only 12 out of 70 studies reported using such methods ([108, 91, 53, 56, 86, 59, 69, 68, 90, 72, 75, 117]). These studies use some form of linear shifting and/or sensor shuffling or mix-up regularization ([169]). An example of linear shifting consists of selecting a different cropping of a trial or segment around the cue in the temporal dimension. The shifted samples end up having the same length but different onsets.
Typically, across all studies in this category, the MEG data were preprocessed before being fed as input to a neural network. This preprocessing phase generally includes down-sampling, band-pass filtering, and either de-noising and/or artifact removal. Of the 70 studies in this category, 12 studies (Decoding: [57, 61, 53, 50] BCI: [91, 97, 100, 98, 92, 93], clinical: [86], event detection: [116]) use discrete wavelet transform to increase signal-to-noise ratio (SNR) in the data as a preprocessing step ([170]). Notably, these 12 studies are spread across only four distinct research groups representing a small portion of the reviewed studies. Five studies employ alternative de-noising techniques ([62, 69, 89, 90, 165]). Additionally, 16 studies use ICA to remove heart and/or eye movement artefacts ([54, 71, 67, 95, 88, 84, 105, 107, 89, 90, 74, 68, 82, 78, 83, 56]). The remaining 38 studies do not explicitly mention de-noising or artifact removal techniques in their preprocessing pipeline.
The studies in this subcategory used MEG data sampled at frequencies between 50 Hz and 2400 Hz. Higher sampling frequencies (e.g., 1000–2400 Hz) were typically used in studies focused on decoding fine-grained temporal dynamics, such as those targeting auditory and motor-related tasks, where capturing rapid neural oscillations (e.g., in the gamma range) is critical. In contrast, lower sampling rates (e.g., 50–250 Hz) were generally the result of preprocessing choices aimed at reducing computational load or were used in studies not reliant on high-frequency information, such as certain clinical applications or feature-based pipelines. Although many studies did not report their original sampling frequency, most MEG systems (e.g., Elekta, CTF, BTi) record at native rates of 1000 Hz or above, suggesting that downsampling is commonly applied during preprocessing. Figure 3 displays the distribution of sampling frequencies across the reviewed studies, with 1000 Hz (26 studies of which 23 being classification studies) and 250 Hz (17 studies of wich 10 are classification studies) being the most frequently used rates, and a median of 600 Hz across all reported values. Still, eight studies omitted details about the sampling frequency, in most cases because the authors extracted features from the MEG data, making the original sampling rates less pertinent ([49, 78, 93, 83, 64, 87, 88, 116]).
3.2.4 Network architectures
ANNs, like any other parametric algorithm, have architectural parameters that play a key role in determining their capacity to learn from data. These parameters —including depth, number of units, and connectivity patterns— affect how well the model can extract and represent relevant features. Optimal configurations vary depending on the type of task and the nature of the input data. In this section, we provide an overview of the most commonly used architectures and their configurations across classification studies.
The most commonly used ANNs for classification-type studies are CNNs. The second most used architecture was multi-layer perceptrons (MLP), followed by recurrent neural network (RNN) architectures. For the CNNs, the depth of the networks varied widely, between one and over 100 hidden layers (up to 164 in one case [97]). This wide range reflects the heterogeneity in input data characteristics, task complexity, and specific CNN architectures employed across the reviewed studies, ranging from relatively shallow custom CNNs to very deep pre-existing architectures like GoogLeNet ([171]) and ResNet ([172]) variants (e.g., [63, 97]). In contrast, the papers that used MLPs essentially limit themselves to a maximum of six hidden layers. This typical depth difference often relates to parameter efficiency: the weight-sharing mechanism in CNNs generally allows for deeper architectures with high-dimensional data such as MEG recordings, whereas the large number of parameters in fully-connected MLP layers can make similar depths computationally costly and difficult to train effectively ([1]).
@@ 3.2.4 Network architectures
Most of the CNN architectures used in the reported MEG classification papers comprise the input layer followed by multiple blocks, typically containing convolution layers followed by activation functions (like rectified linear unit, ReLU) and often pooling layers (like max pooling) for downsampling. At the end, one or two fully connected layers are followed by an activation function. For the convolution layers, the critical parameters are filter size, stride, padding, and number of channels. For fully connected layers, the crucial parameter is the number of units, also called the width of the layer.
Although an in-depth discussion of parameter initialization techniques is beyond the intended scope of this review, it is important to keep in mind that initialization methods (e.g., random, Xavier, or He initialization) are important and can profoundly influence model performance and stability (see [173]). Similarly, the choice of activation function (e.g., Sigmoid, Tanh, ReLU and its variants) applied after convolutional or fully connected layers is another critical design decision impacting network behavior and learning dynamics.
3.2.5 Training and validation techniques
@@
Beyond the architecture used and data (pre-)processing, details about training parameters, loss function, optimizer, and regularization techniques are essential to replicate the results of such studies. While training parameters may not have as significant an impact as architecture parameters, they remain essential and can heavily impact training time and performance. Out of the 70 classification studies, 17 do not mention training parameters. The remaining studies generally mention learning rate, batch size, loss function, optimization method, and regularization techniques. Batch sizes, regularization techniques (dropout, L1 and L2 norm weight usage), number of epochs, early stop criteria, learning rates, and momentum parameters are crucial to watch out for. While the specific training parameters used in each study are crucial for reproducibility, the high degree of variability across different experimental settings and goals makes a detailed summary within this review text impractical. Instead, to provide a resource for readers interested in specific implementations, further details on the reported parameters for individual studies can be found in the comprehensive online table (https://tinyurl.com/ub3s5mr).
@@
Of the studies that specify the optimization algorithm utilized for training their network, 29 employed adaptive moment estimation (Adam [174]) or one of its variants, while four used SGD. Adam, uses both the first-order and second-order moments of the gradient to adapt the learning rate, whereas SGD only uses the first-order moment. Adam is known for its faster convergence in some architectures ([175]), but it can quickly plateau after convergence ([176]) and may perform worse than SGD in specific use cases ([145]).
@@
Validation techniques are essential to training any model in a classification setting, enabling the proper evaluation of the model’s performance. In a setting where generalizing to new subjects is essential, such as BCI or CAD, it is crucial to exclude multiple subjects’ data from the training process. This ensures the test of the trained algorithm’s generalization capabilities on entirely new and unseen data. It is also essential to evaluate and fine-tune the algorithm’s performances on training and validation datasets that do not share information with each other or the test set. Among the 71 studies in this category, 26 used the K-fold validation technique (K varying from two to ten), 29 used a simple train/valid/test set split, 11 used leave one out or leave P out at the subject level. Five did not mention any validation technique in their analysis pipeline. The choice between a simple train/valid/test split and cross-validation methods (like K-fold or leave-subject-out) often depends on dataset size and computational resources. While cross-validation provides more robust performance estimates and is recommended when data is scarce, as it leads to a less noisy estimate of performance ([177]), simple splits are frequently used in deep learning when dealing with very large datasets where the computational cost of K-fold is high and a single split is deemed sufficient for stable evaluation. Reflecting this trade-off, the studies in our review using simple splits often involved substantially larger datasets compared to those employing cross-validation (see tables 2 and 3). The leave one subject out (LOSO) cross-validation strategy is particularly relevant when subject generalization is essential and the amount of available data is small. Furthermore, when using K-fold or random splits, it is crucial to consider potential subject data leakage between folds or sets, ensuring that the evaluation reflects true generalization rather than subject identification.
@@
ANNs are powerful and can tackle a large set of problems; however, they present challenges in terms of interpretation, implementation, and tuning. In some instances, it is unnecessary to use ANNs for problems that simpler algorithms can solve. This is why it is essential when using ANNs to check whether we are using the right tool for the right problem by computing a baseline of performances that can be reached with more classical machine-learning approaches. Unlike ANNs, which can implicitly learn features, classical algorithms require careful feature selection to effectively manage the high dimensionality of MEG data. Among the papers in the ’Classification’ category, 27 out of the 70 studies did not compare their ANN’s performance to simpler algorithms’ performances.
@@
A critical part of decoding studies is understanding what aspect of the data allowed decoding. It is essential to find where and how information is encoded in the data to understand more about the brain. When using ML algorithms, especially ANNs, it can be hard to make sense of what is happening in the latent space. This is why using visualization tools will help with the interpretability of the network. Of the 70 papers in the ’Classification’ category, only 16 included interpretation or visualization techniques applied to the network. These techniques varied widely, including methods such as visualizing activation patterns or feature maps ([178]), generating feature importance or contribution maps (e.g., using permutation feature importance [179], saliency methods [180], or additive feature attribution approaches [181]), analyzing performance across time or frequency, examining network-derived connectivity, and applying specific tools like gradient-weighted class activation mapping (Grad-CAM [182]). This list provides illustrative examples but is not exhaustive. Further details on methods used in specific studies can be found in the supplementary online table (https://tinyurl.com/ub3s5mr). Only one of these studies was a BCI sub-categorized study ([95]), two were event detection studies ([105, 114]), four were clinical sub-categorized studies ([82, 78, 86, 88]), and the remaining 9 were decoding studies ([51, 55, 54, 69, 56, 67, 68, 73, 74]).
3.2.6 Imbalanced datasets
Despite the growing interest in using ANNs for MEG-based classification, relatively few studies explicitly address the challenge posed by imbalanced datasets. This issue is especially prevalent in clinical and event detection studies, where one class—such as a rare neurological condition or the presence of epileptic spikes—may be underrepresented in the training data. Class imbalance can significantly bias the model’s performance, as neural networks tend to favor the majority class unless corrective measures are implemented.
Among the 70 classification studies reviewed, only a handful explicitly mention strategies to deal with class imbalance. When addressed, this was most commonly achieved through the use of weighted loss functions, such as weighted cross-entropy, which assign greater penalty to errors on the minority class. Some studies adopted data-level strategies such as upsampling the minority class or synthetic data generation through techniques like synthetic minority oversampling technique (SMOTE [183]), although these methods were rarely discussed in detail. Others relied on evaluation metrics that are more robust to imbalance, such as the F1 score, precision-recall curves, or area under the ROC curve (AUC), rather than classification accuracy. In general, however, most studies did not report class distributions or provide a rationale for their evaluation choices, making it difficult to assess the impact of imbalance on their findings.
The limited attention given to this issue is concerning, given the known sensitivity of neural networks to unbalanced data, particularly when applied to high-dimensional signals like MEG. Future work in this area would benefit from more systematic reporting of class distributions, explicit justification of performance metrics, and comparative evaluations of different strategies for mitigating imbalance ([184]). Doing so would improve the interpretability and reproducibility of results and allow for more meaningful comparisons across studies.